Adv. Math. Econ. 8, 397-407 (2006)

Solution-based congestion games

Dov Monderer

Fculty of Idustorial Engineering and Management, Technion-Israel Institute of Technology, Haifa,Israel Received: September 7, 2005 Revised: September 23, 2005 JEL classification: C71, C72 Mathematical Subject Classification (2000): 91A10, 91A12 Abstract. In this paper we develop the theory of potential of cooperative games for semivalues, characterize congestion models that are defined by semivalues, and suggest an application of these results to combinatorial auctions, which may explain the success of the Iowa electronic market.